OpenStudy (tomektomeknyc):

If log 94.3 = 1.9745 then log 0.0943 = a) 1.9745- 2 b) 10.9745- 2 c) 8.9745 d) - 2.9745

OpenStudy (amoodarya):

log 0.0943 =log 94.3 * 0.001 so log (94.3) +log(0.001) =log(94.3) +log(10^(-3)) =log94.3 -3

OpenStudy (amoodarya):

-3+log 94.3 =-3+1.9745

OpenStudy (tomektomeknyc):

Thank you very much. So what is the answer? I guess none of answers provided is correct?